OpenStudy (anonymous):

Need help! Δ DOG has coordinates D (3, 2), O (2, -4) and G (-1, -1). A translation maps point D to D' (2, 4). Find the coordinates of O' and G' under this translation.

OpenStudy (anonymous):

O' (-1, 1);G' (1, -2) O' (3, 0);G' (-1, -1) O' (0, -3);G' (-3, 0) O' (1, -2);G' (-2, 1)

OpenStudy (anonymous):

@akshay1234 @Chris911 @beccaboo333 @halorazer @frrrrrrr @FunMedley @

OpenStudy (anonymous):

@Whitemonsterbunny17 @doulikepiecauseidont @FreyaBeauchamp @Cydney_Morgan @Vitter @bbcream14 @jamielovesmickey @jzellem1004 @jim_thompson5910 @kryptonite1218

OpenStudy (anonymous):

It's kind of against the code of conduct to tag so many people.

OpenStudy (anonymous):

lol I know some of the people and I am just trying to get help on this question for my assignment @halorazer

OpenStudy (anonymous):

Since it's a translation, you need to add/subtract fixed values from the x and y coordinates of each point. Point D(3, 2) is translated to D'(2, 4). Clearly we subtract 1 from the x coordinate and add 2 to the y coordinate. Likewise the translated points of G and O will be O(2-1, -4+2) = O(1, -2) G(-1-1, -1+2) = G(-2, 1) So option D is the right answer.